OpenStudy (jjuden):

help Factor out the greatest common factor. 4x + 36x^2

OpenStudy (skullpatrol):

$$\Huge 4*?=36$$

OpenStudy (jjuden):

9

OpenStudy (skullpatrol):

$$\Huge 4x + 36x^2=4x + 4*9x^2$$ right?

OpenStudy (jjuden):

yes

OpenStudy (skullpatrol):

What do the terms $$4x$$ and $$4*9x^2$$ have in common, that is what do the BOTH share?

OpenStudy (jjuden):

4

OpenStudy (skullpatrol):

Yes, what else?

OpenStudy (jjuden):

x

OpenStudy (skullpatrol):

Yes, therefore the GCF is 4x

OpenStudy (skullpatrol):

$$\Huge 4x + 36x^2=4x(1+9x)$$

OpenStudy (skullpatrol):

$$\Huge ab+ac=a(b+c)$$

OpenStudy (jjuden):

ok